The temperature of a gas in a rigid container is decreased from 127.00°c to 27.0°c. the intitial pressure was 1000mm Hg. what is the new pressure in mm Hg?
1
Expert's answer
2020-04-15T13:58:59-0400

From ideal gas law p*V = n*R*T under the condition of constant V (rigid container) and constant n (quantity of moles of gas), it follows that:

p1/T1 = p2/T2

where T is in Kelvin (°c + 273). Thus, the new pressure is:

p = 1000/(127+273)*(27+273) = 750 mm Hg